Find NOWC (Net Operating Profit After Taxes) for 2013

Asset Year

2014

2013

2012

Cash

14,000

7,282

9,000

Acct Receivable

878,000

632,160

351,200

Inventories

1,716,480

1,278,360

715,200

FA

817,040

939,790

344,800

Acct Payable

436,800

524,160

145,600

Accrual

408,000

489,600

136,000

Income Statement Year

2014

2013

2012

EBIT

510,640

-690,560

209,100

TAX

40%

$ 2,580,720
$ 1,843,832
- $ 414,336
$ 306,384
Not enough information
Answer #1

Net operating profit after tax (NOPAT) = EBIT(1-t) = -690560 (1-0.40) = -414336

Answer : - $414336 (THUMBS UP PLEASE)